OpenStudy (samanthagreer):

Alicia estimated her test score to be 85. Her actual score was 80. What was the percent error in Alicia’s estimate? Round your answer to the nearest tenth of a percent. %

OpenStudy (anonymous):

what

OpenStudy (samanthagreer):

@TheEdwardsFamily

OpenStudy (anonymous):

(85-80)/80=5/80=1/16=0.0625=6.25%

OpenStudy (anonymous):

rounded 6%

OpenStudy (anonymous):

@Allenismyname Please erase comment and give the learner tips of the question.

OpenStudy (anonymous):

Do you understand? samanthagreer

OpenStudy (samanthagreer):

well no cause she did put how she got that answer and so it is 6% correct